Lon Timur is an accounting major at a midwestern state university located approximately 60 miles from a major city. Many of the students attending the university are from the metropolitan area and visit their homes regularly on the weekends. Lon, an entrepreneur at heart, realizes that few good commuting alternatives are available for students doing weekend travel. He believes that a weekend commuting service could be organized and run profitably from several suburban and downtown shopping mall locations. Lon has gathered the following investment information.

1. Five used vans would cost a total of $75,096 to purchase and would have a 3-year useful life with negligible salvage value. Lon plans to use straight-line depreciation.
2. Ten drivers would have to be employed at a total payroll expense of $48,600.
3. Other annual out-of-pocket expenses associated with running the commuter service would include Gasoline $15,800, Maintenance $3,100, Repairs $3,900, Insurance $4,300, and Advertising $2,500.
4. Lon has visited several financial institutions to discuss funding. The best interest rate he has been able to negotiate is 15%. Use this rate for cost of capital.
5. Lon expects each van to make ten round trips weekly and carry an average of six students each trip. The service is expected to operate 30 weeks each year, and each student will be charged $12 for a round-trip ticket.


Click here to view PV table.

(a)

Determine the annual (1) net income and (2) net annual cash flows for the commuter service. (Round answers to 0 decimal places, e.g. 125.)

Net income $

Net annual cash flows $


(b)

Compute (1) the cash payback period and (2) the annual rate of return. (Round answers to 2 decimal places, e.g. 10.50.)

Cash payback period

years
Annual rate of return

%


(c)

Compute the net present value of the commuter service. (Round answer to 0 decimal places, e.g. 125. If the net present value is negative, use either a negative sign preceding the number eg -45 or parentheses eg (45). For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

Net present value
Answer #1

(a)

Net Income=Sales- Depreciation - Payroll expense- Out of pocket expenses

Sales = No of vans X No of Trips per week X No of weeks per year X No of Students per van X Price for each trip

=5 X 10 X 30 X 6 X12

= $108000

Payroll Expense = $48600

Out of Pocket Expense=$ 15800+ $ 3100+ $ 3900 +$ 4300+ $ 2500

=$ 29600

Depreciation =Total Cost / No of useful life

=75 ,096/ 3

=$ 25,032

Net Income= $108000-48600-29600-25032   

= $ 4768

Net annual cash flows = Sales- Out of pocket expenses - Payroll Cost

=108000-48600-29600

=$ 29,800

( b)

Cash back period = Initial Investment / Annual cash flows

Cash Back period =75096/29800

= 2.52 years

Annual Rate of return = Annual income/ Initial investment *100

= 4768/75096*100

=6.35 %

(c) Net present Value = PV of cash flows - Initial investment

P V Of Cash Flows - Pv values taken from chart with cost of capital 15%

PV = 29800* 0.8697+ 29800*.75614+29800*.65752

=25917+22532+19594

= $ 68043

Net present Value = $ 68043- $ 75096

=$ 7053